OpenStudy (anonymous):

pleaseeee helpppp me understand thiss... The polynomial x 3 + 5x 2 -57x -189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions? ( Hint: The height is greater than the depth.) A. height: 19 in. depth: 5 in. B. height: 21 in. depth: 5 in. C. height: 19 in. depth: 7 in. D. height: 21 in. depth: 7 in.

OpenStudy (amistre64):

divide out the (x+3) from the givin poly and factor the results

OpenStudy (amistre64):

or we can work it more starightforward (x+3) (ax^2 + bx + c) has to equate to the given poly multiply it out and compare parts

OpenStudy (amistre64):

or, leave it in all factored form, and then multiply (x+3)(x-h)(x-d) and compare parts .... there are a few approaches to it, but these seem to be the most simplest to comprehend

OpenStudy (amistre64):

(x+3) (ax^2 + bx + c) x(ax^2 + bx + c) +3(ax^2 + bx + c) ax^3 + bx^2 + cx +3ax^2 + 3bx + 3c ax^3 + (b+3a)x^2 + (c+3b) x + 3c 1x^3 + 5 x^2 -57 x -189 comparing parts a=1 c=-189/3 = -63 b+3a = 5, b = 2, since a=1 so our quadratic is: x^2 +2x -63 does it factor?

OpenStudy (amistre64):

9 and -7 are -63, and add to 2 (x+3)(x-9)(x+7) defines the parts, what are the parts when x=12?

OpenStudy (amistre64):

+9 an -7 ....

OpenStudy (anonymous):

im sorry im still really confused

OpenStudy (amistre64):

the roots are the opposite of the dohickies ... got a little confused at the end lol

OpenStudy (amistre64):

you have to find someway to determine the quadratic that (x+3) multiplies to to get the stated cubic .. right?

OpenStudy (anonymous):

okay right, i get it but im just confused about what to do at the end.